# Native Design Bias: Studying the Impact of English Nativeness on Language Model Performance

### v2

@@ 7 Conclusion
In this work, we analyze bias in LLMs towards native English speakers. We analyze whether models perform better for native compared to non-native English speakers and whether the models are even further tuned towards Western native English speakers. We find that there are significant performance differences between native and non-native prompts. More specifically, models become more inaccurate for the latter and this effect is even more pronounced for the Western native English versus not Western native English comparison for objective classification tasks. Furthermore, we find a strong anchoring effect when information about the user’s nativeness is added or inferred by the model. For our experiments, we used a newly collected dataset consisting of over 12,000 unique prompts from a diverse set of annotators.
@@ 8 Limitations
Our dataset contained a very diverse set of annotators. Nevertheless, it would be interesting to have more study participants for every sub-population, such that general findings at sub-population level could be made as well. Furthermore, our experiments contained mostly annotators having a self-reported level of English of C1 and C2. It would be very interesting to analyze the effects on the performance of LLMs when prompted by people having different levels of English as this will probably also be impactful. Additionally, our results were only gathered for five different models. It would be insightful to extend this analysis to more models, as every model is trained differently and therefore these design choices might lead to different biases within the model. An important limitation of using LLMs and especially the closed-source variant thereof, is the lack of reproducibility of the results. We make available a multilingual dataset, however, have only analyzed the English answers. We leave the analysis of bias in the multilingual dataset for future research. Finally, we acknowledge that our way of evaluating, using the LLM-generated answers as reference sentences for the generative sentences, might be affected by the prompt of the original prompt writer. We have also analyzed the performance when letting the model first annotate the example and then use this to gather the response from the chatbot. However, we find that this resulted in large performance drops for the smaller models. Moreover, for the models that did offer adequate results, we find similar results as when we used the original prompt. The results are included in Appendix N.
@@ 9 Ethical considerations
Acknowledgments
This research was funded by the Statistics Flanders research cooperation agreement on Data Science for Official Statistics.
@@ Appendix A Dataset overview
We used the datasets as they were assembled by Mishra et al. (2022) and Wang et al. (2022). Table 6 shows an overview of the selected datasets, together with their task ID in the original instructions dataset. The task definition given in the table is the one we used when prompting the models. For CNN Dailymail and CODA19, this differs from the original task definition in the dataset because we flipped the task. Instead of letting our annotators write the article, we asked them to write the summary or title respectively. Datasets Abductivenli, Timetravel, Amazonfood, McTaco, TweetQA, and Commonsense are thus classification tasks, while datasets StoryCloze, CNN Dailymail, CODA19, and Paraphrase are generation tasks.
@@
task069 Abductivenli In this task, you will be shown a short story with a beginning, two potential middles, and an ending. Your job is to choose the middle statement that makes the story coherent / plausible by writing 1̈ör 2̈ïn the output. If both sentences are plausible, pick the one that makes most sense.
@@
task020 Mctaco The answer will be ’yes’ if the provided sentence contains an explicit mention that answers the given question. Otherwise, the answer should be ’no’. Instances where the answer is implied from the sentence using ïnstinctör c̈ommon sense(̈as opposed to being written explicitly in the sentence) should be labeled as ’no’.
task241 TweetQA In this task, you are given a context tweet, a question and the corresponding answer of the given question. Your task is to classify this question-answer pair into two categories: (1) ÿesïf the given answer is right for question, and (2) n̈oïf the given answer is wrong for question.
@@
task295 Commonsense In this task, you are given an impractical statement. You are also given three reasons (associated with Ä,̈ B̈,̈ C̈)̈ explaining why this statement doesn’t make sense. You must choose the most corresponding reason explaining why this statement doesn’t make sense.
@@ Appendix D Dataset Statistics -Annotations
Native language Number of annotators Languages Validation rate
@@
Set ids Native or not Western native or not Total
Native Non-native Western Not Western
@@ D.1 Prompt length
Dataset ids Native or not Western native or not
native non-native not Western native Western native
@@ D.2 Time analysis annotators
Set-ids Native or not Western native or not
native non-native Western native not Western native
@@ Appendix E Prompt templates
Guess Native
system prompt: You are a helpful assistant.
chat history1:
User: Guess whether the writer of the following prompt is a native or non-native English speaker: [ANNOTATOR PROMPT],
chat history2:
User: Guess whether the writer of the following prompt is a native or non-native English speaker: [ANNOTATOR PROMPT],
Assistant:[RESPONSE],
User: Next, execute the following task taking this information into account. [task definition],
Assistant:Understood,
User: [Full English Prompt]
@@
Appendix F Checkpoints models
@@
Haiku was made by Anthropic (Anthropic, 2024). We used claude-3-Haiku-20240307.
Sonnet was made by Anthropic (Anthropic, 2024). We used claude-3-Sonnet-20240229.
Qwen 7B is an open source model made by the Alibaba group (Bai et al., 2023). We used Qwen/Qwen1.5-7B-Chat
@@ Appendix G Evaluation metrics
BLEU. [^10] https://www.nltk.org/api/nltk.translate.bleu_score.html This established performance measure relies on counting the n-gram overlap between candidate and reference sentence(Papineni et al., 2002).
ROUGE. [^11] https://huggingface.co/spaces/evaluate-metric/rouge This metric is commonly used for summarization tasks and calculates the n-gram recall between a reference and candidate summary. We employ ROUGE-1, ROUGE-2, and ROUGE-L (Lin, 2004).
BERT-score. [^12] https://huggingface.co/spaces/evaluate-metric/bertscore This evaluation metric calculates and aggregates a similarity score between the BERT input embeddings for both the candidate and reference sentences. We employed microsoft/deberta-xlarge-mnli which was mentioned to be the best-performing model at the moment of writing the paper (Zhang et al., 2019).
BART-score.131313github.com/neulab/BARTScore This evaluation measure approaches the evaluation as a text-generation task, including all parameters learned during the pre-training phase (Yuan et al., 2021).
All metrics were implemented using the default parameters of the cited packages.
@@ Appendix H Distribution Amazon food reviews
Figures 5 and 6 show an overview of the wrong predictions of the AmazonFood review dataset for the different groups and models for one of the three runs. This shows the distribution between what was predicted and what should be predicted. We only consider here the cases where the model predicted one of the given ratings, and excluded cases where no prediction was given. As shown, for both the native and Western native group, we find a large amount of misclassification for the highest rating. Additionally, neutral is not often predicted for these classes compared to the other groups.
(a) Overview of the wrong predictions for the native English speakers.
(b) Overview of the wrong predictions for the non-native English speakers.
Figure 5: Overall misclassification for native and non-native English speakers
(a) Overview of the wrong predictions for the Western native English speakers.
(b) Overview of the wrong predictions for the not Western native English speakers.
Figure 6: Overall misclassification for Western native English speakers and not Western native English speakers

### v3

@@ 7 Conclusion
In this work, we analyze bias in LLMs towards native English speakers. We analyze if models perform better for native compared to non-native English speakers and whether the models are even further tuned towards Western native English speakers. We find that there are performance differences between native and non-native prompts. More specifically, models are most accurate for the Western-native English speakers on objective classification tasks. A slightly lower performance is shown for the NWN group compared to the NN, nevertheless, we show that this is mostly model-dependent. Both GPT models seem even to prefer NN over NWN, while the other models in our analysis show similar performance for both native groups. Furthermore, we find a strong anchoring effect when information about the user’s nativeness is added for objective classification tasks. Generative tasks seem to be in general more robust against this native bias, probably due to the longer context length and the optimization of the used models towards these generative tasks. For our experiments, we used a newly collected dataset consisting of over 12,000 unique prompts from a diverse set of annotators.
@@ 8 Limitations
Our dataset contained a very diverse set of annotators. Nevertheless, it would be interesting to have more study participants for every sub-population, such that general findings at sub-population level could be made as well. The annotations in the dataset are done by annotators from different groups. However, there is an imbalance in number of native English speakers compared to the number of non-native English speakers. Furthermore, our experiments contained mostly annotators having a self-reported level of English of C1 and C2. It would be very interesting to analyze the effects on the performance of LLMs when prompted by people having different levels of English as this will probably also be impactful. Additionally, our results were only gathered for five different models. It would be insightful to extend this analysis to more models, as every model is trained differently and therefore these design choices might lead to different biases within the model. An important limitation of using LLMs and especially the closed-source variant thereof, is the lack of reproducibility of the results. We make available a multilingual dataset, however, have only analyzed the English answers. We leave the analysis of bias in the multilingual dataset for future research. Additionally, some of the datasets contain a Western focus in terms of the topics that are discussed (CNN Dailymail, TweetQA, and McTaco). While other datasets, like the Amazon Food reviews, are based on user-generated content and may be less culturally specific, we recognize that the overall selection may still reflect Western contexts. Finally, we acknowledge how the LLM-as-a-judge implementation for gathering generative results might be suboptimal to human annotators due to model-specific biases. Therefore, we chose a different LLM than the ones we will evaluate to serve as a judge to avoid self-preference bias and we manually validated a sample. To further assess the reliability we have included both the correlation and cohen’s kappa score. Given the high score for both metrics between the manual annotations and the LLM annotations, we assume that the LLM annotations are representative.
@@ Appendix A Dataset overview
We used the datasets as they were assembled by naturalinstructions and supernaturalinstructions. Table 1 shows an overview of the selected datasets, together with their task ID in the original instructions dataset. The task definition given in the table is the one we used when prompting the models. For CNN Dailymail and CODA19, this differs from the original task definition in the dataset because we flipped the task. Instead of letting our annotators write the article, we asked them to write the summary or title respectively. Datasets Abductivenli, Timetravel, Amazonfood, McTaco, TweetQA, and Commonsense are thus classification tasks, while datasets StoryCloze, CNN Dailymail, CODA19, and Paraphrase are generation tasks.
@@
task069 Abductivenli In this task, you will be shown a short story with a beginning, two potential middles, and an ending. Your job is to choose the middle statement that makes the story coherent / plausible by writing "1" or "2" in the output. If both sentences are plausible, pick the one that makes most sense.
@@
task020 Mctaco The answer will be ’yes’ if the provided sentence contains an explicit mention that answers the given question. Otherwise, the answer should be ’no’. Instances where the answer is implied from the sentence using "instinct" or "common sense" (as opposed to being written explicitly in the sentence) should be labeled as ’no’.
task241 TweetQA In this task, you are given a context tweet, a question and the corresponding answer of the given question. Your task is to classify this question-answer pair into two categories: (1) "yes" if the given answer is right for question, and (2) "no" if the given answer is wrong for question.
@@
task295 Commonsense In this task, you are given an impractical statement. You are also given three reasons (associated with "A", "B", "C") explaining why this statement doesn’t make sense. You must choose the most corresponding reason explaining why this statement doesn’t make sense.
@@ Appendix D Dataset Statistics -Annotations
Native
language
Number of
annotators
Languages Validation rate
@@
Set ids
Native or not
Western native or not
Total
Native
Non-native
Western
Not Western
@@
Furthermore, the annotators have reported their level of English proficiency and the frequency of which English was spoken. We provide this information for the non-native speakers in Tables 6 and 7.
English proficiency level
Number of non-native
annotators
C2 31
C1 41
B2 13
B1 1
Table 6: Overview of the self-reported English proficiency of the non-native annotators.
English usage frequency
Number of non-native
annotators
Daily 60
A few times per week 21
Once a week 4
Less than once a week 1
Table 7: Overview of the self-reported frequency of English usage of the non-native annotators.
@@ D.1 Prompt length
Dataset ids
Native or not
Western native or not
native
non-native
not Western native
Western native
@@ D.2 Time analysis annotators
Set-ids Native or not
Western native or not
native non-native
Western native
not Western native
@@ Appendix E Prompt templates
Appendix F Checkpoints models and hyperparameters
@@
Haiku was made by Anthropic (anthropic2024claude). We used claude-3-Haiku-20240307.
Sonnet was made by Anthropic (anthropic2024claude). We used claude-3-Sonnet-20240229.
Qwen 7B is an open source model made by the Alibaba group (bai2023qwen). We used Qwen/Qwen1.5-7B-Chat
We have generated the responses for the classification tasks with a temperature set at 0 and for the generation tasks at 0.7
@@ Appendix G Evaluation metrics
All generative performance metrics were gathered using Llama-3.3-70B-Instruct. Below, the prompts for the different metrics were given. These are based on the ones used for the Summeval dataset by Bavaresco2024JUDGE_BENCH. Depending on the generative task, the prompts differ slightly. we use the following structure: Final prompt= ""start_prompt + doc_gen: [model generated reply] + metric"" where the start prompt for fluency is:
’You will be given a doc_gen generated based on a doc_comp. Your task is to rate the doc_gen on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
and the start prompt for the other metrics is: You will be given a doc_gen and a doc_comp. Your task is to rate the doc_gen on one metric. Please make sure you read and understand these instructions carefully. Please keep this document open while reviewing, and refer to it as needed.
doc_gen and doc_comp depend on the dataset. They are a ’closing sentence’ and a ’story’ for Story Cloze; a ’news article’ and a ’summary’ for CNN Dailymail; an ’article’ and a ’title’ for CODA19; and a ’paraphrased sentence’ and a ’sentence’ for Paraphrase.
The metric is one of the following:
Fluency:
Evaluation Criteria: Fluency (1-3): the quality of the doc_gen in terms of grammar, spelling, punctuation, word choice, and sentence structure. Assign a score on a scale of 1 to 3 where: - 1: Poor. The doc_gen has many errors that make it hard to understand or sound unnatural.- 2: Fair. The doc_gen has some errors that affect the clarity or smoothness of the text, but the main points are still comprehensible. - 3: Good. The doc_gen has few or no errors and is easy to read and follow. Evaluation Form (scores ONLY):- Fluency:
Coherence.
The definition depends on the dataset. For CNN Dailynews and CODA19 it is as follows:
Evaluation Criteria: Coherence (1-5) - the collective quality of all sentences. We align this dimension with the DUC quality question of structure and coherence whereby the doc_gen should be well-structured and well-organized. The doc_gen should not just be a heap of related information, but should build from sentence to a coherent body of information about a topic. Evaluation Steps: 1. Read the doc_comp carefully and identify the main topic and key points.2. Read the doc_gen and compare it to the doc_comp. Check if the doc_gen covers the main topic and key points of the doc_gen, and if it presents them in a clear and logical order.3. Assign a score for coherence on a scale of 1 to 5, where 1: Very low coherence ; 2: Low coherence; 3: Mediocre coherence ; 4: High coherence ; 5: Very high coherence. Evaluation Form (scores ONLY):- Coherence:
For Paraphrase it is as follows:
Evaluation Criteria: Coherence (1-5) - The overall quality of the paraphrased sentence in terms of logical flow, structure, and alignment with the original sentence. A coherent paraphrase should preserve the meaning of the original sentence, avoid redundancy, and introduce variation without altering the main idea. The paraphrased sentence should not feel disjointed or incomplete but should read smoothly as a standalone sentence. Evaluation Steps: 1. Read the doc_comp carefully and identify the main topic and key points. 2. Read the doc_gen and compare it to the doc_comp. 3. Assign a score for coherence on a scale of 1 to 5, where 1: Very low coherence ; 2: Low coherence; 3: Mediocre coherence ; 4: High coherence ; 5: Very high coherence. Evaluation Form (scores ONLY): - Coherence:
For Story Cloze it is as follows: Evaluation Criteria: Coherence (1-5) - the collective quality of all sentences. We align this dimension with the DUC quality question of structure and coherence whereby the sentences should be well-structured and well-organized. The sentences should not just be a heap of related information, but should build from sentence to a coherent story.Evaluation Steps: 1. Read the doc_comp carefully and identify the main topic and key points. 2. Read the doc_gen and compare it to the doc_comp. Check if the sentences are clear and in a logical order. 3. Assign a score for coherence on a scale of 1 to 5, where 1: Very low coherence ; 2: Low coherence; 3: Mediocre coherence ; 4: High coherence ; 5: Very high coherence. Evaluation Form (scores ONLY): - Coherence:
Relevance.
The definition depends on the dataset. For Story Cloze it is as follows:
Evaluation Criteria: Relevance (1-5) - The degree to which the generated doc_gen effectively reflects the main themes and purpose of the doc_comp. A relevant closing sentence should provide a meaningful and appropriate conclusion, aligning with the tone and key points of the narrative. Evaluation Steps: 1. Read the doc_comp and the doc_gen carefully. 2. Compare the doc_gen to the doc_comp and identify the main points of the doc_comp. 3. Assess how well the doc_gen concludes the doc_comp, and how much irrelevant or redundant information it contains. 4. Assign a relevance score from 1 to 5 where 1: Very low relevance ; 2: Low relevance; 3: Mediocre relevance ; 4: High relevance ; 5: Very high relevance. Evaluation Form (scores ONLY): - Relevance:
For all other datasets it is as follows:
Evaluation Criteria: Relevance (1-5) - inclusion of important content from the doc_comp. The doc_gen should include all important information from the doc_comp. Evaluation Steps: 1. Read the doc_comp and the doc_gen carefully. 2. Compare the doc_gen to the doc_comp and identify the main points of the doc_comp. 3. Assess how well the doc_gen covers the main points of the doc_comp, and how much irrelevant or redundant information it contains. 4. Assign a relevance score from 1 to 5 where 1: Very low relevance ; 2: Low relevance; 3: Mediocre relevance ; 4: High relevance ; 5: Very high relevance. Evaluation Form (scores ONLY): - Relevance:
@@ Appendix H Distribution Amazon food reviews
Figure 10 shows an overview of the wrong predictions of the AmazonFood review dataset for the different groups and models for one of the three runs. This shows the distribution between what was predicted and what should be predicted. We only consider here the cases where the model predicted one of the given ratings, and excluded cases where no prediction was given. As shown, for both the native and Western native group, we find a large amount of misclassification for the highest rating. Additionally, neutral is not often predicted for these classes compared to the other groups.
(a) Overview of the predictions for the Western native English speakers.
(b) Overview of the predictions for the non-native English speakers.
(c) Overview of the predictions for the native English speakers that are not western native.
Figure 10: Overall classifications for Western native, native that are not Western native, and non-native English speakers